OpenStudy (anonymous):

Splott Fizzle 2x + 4 = –3x + 14 how do I solve it

OpenStudy (akashdeepdeb):

WELCOME TO OPENSTUDY!! :D

OpenStudy (akashdeepdeb):

2x + 4 = –3x + 14 2x+3x = +14-4 now try.

OpenStudy (anonymous):

5x=10 5/5x=10/5 x=2